the width of a rectangular room is four less than the length and its perimeter is greater than 28. what are the least possible integral measures of the length and width

Answer by rothauserc(4718)

let w be width and l be length, then
:
w = l - 4
:
Perimeter = 2w + 2l
:
2(l-4) + 2l > 28
:
2l -8 + 2l > 28
:
4l > 36
:
l > 9
:
l = 9 + 1 = 10
:
w = 10 - 4 = 6
:
*******************************
length is 10 and width is 6
:
check the answer
:
2(10) + 2(6) > 28
:
32 > 28
:
our answer checks
********************************
: